QUESTION FIFTEEN


On Monday the 1" of November 2004 the dollar/euro exchange rate was 1.2748 and the dollar/euro 12


months forward exchange rate was 1.2717.


a. Calculate the annual forward premium of the euro and the dollar.


b. Explain whether the forward premium indicates that investors expect the curo to appreciate or


depreciate in the future.


c. If you expect that the euro will appreciate in the future, should you buy or sell a forward euro


contract? [15 MARKS]

1
Expert's answer
2022-01-03T10:56:48-0500

"Solution"

A)

"Forward\ premium(discount)=\frac{forward-spot \ rate }{spot \ rate} × \frac{12}{no.\ of\ months }"


"Fw \ P(d)=\frac{1.2717-1.2748 }{1.2748}\times \frac{12}{12}\\
=-0.024\\
=(2.4\%)"

It's being sold at a discount.

B) they expect the euro to depreciate as investors will sell euros on the spot to buy the dollar.

C) buy the euro and sell the dollar in forward

The forward rate will go up , the spot rate will go down so the euro Forward will be sold at a premium.